Answers: AB = -15.3 AB-T-15.3 A4 - 49.25 km in KN T42.42 GH = 149.25 knDetermine the force in each member of the loaded truss. All triangles are 3-4-5. Enter a positive number if the member is in
Answer #1

couldet Here all Huangles are 25NN 22XN J.V-5 patan 72. ) - 36.589 - Silid A T PIN consider fx=0 Mx + 25 cos 53.15 +22 - 0 THFFD coneides joint F. {Fz=0 E Fx=0 227 FD 20336.86 FCB los 36,86=0 Fed=0 FGF NU 2 Focos 36.86=-22 TFCD = -13-746N (0)] FcR =